Hello, I am a junior struggling with her summer homework.
Here is the question:
An arc of a circle has a measure of 80 degrees and a length of 88pi centemeters. What is the diameter of the circle?
Please help!
Thanks a lot!

Answer by Alan3354(8321)

An arc of a circle has a measure of 80 degrees and a length of 88pi centemeters. What is the diameter of the circle?
----------------
The entire circle, 360 degs, has a circumference of 2pi*r
80 degs is 80/360 of the circumference, = 2/9.
So, 2pi*r*2/9 = 88pi
4pi*r/9 = 88pi
r/9 = 22
r = 198 cm
dia = 396 cm